Write an equation of a line parallel to y = 2 over 3x −5 in slope-intercept form that passes through the point (−6, −1).
Answer by lwsshak3(11628)

Write an equation of a line parallel to y = 2 over 3x −5 in slope-intercept form that passes through the point (−6, −1).
***
y=2/3x-5
slope=2/3
..
line thru point (-6,-1)
y=2/3x+b
solve for b:
-1=(2*-6)/3+b
b=3
equation: y=(2/3)x+3